OpenStudy (anonymous):

Find the slope of the tangent using f(x+h) - f(x) / h y = 8 /x where x = 2

OpenStudy (bahrom7893):

y = 8/x, find the derivative using the limit definition..

OpenStudy (bahrom7893):

all of this will be lim as x->0 (f(x+h)-f(x))/h = (2/(x+h) - 8/x)/h

OpenStudy (anonymous):

yea i did..but i got 0 on the nominator

OpenStudy (anonymous):

isnt it 8 / x + h - ( 8/x) / h?

OpenStudy (bahrom7893):